# Zero-Knowledge Proofs in Sublinear Space

### v1

@@ Abstract
Modern zero-knowledge proof (ZKP) systems, essential for privacy and verifiable computation, suffer from a fundamental limitation: the prover typically uses memory that scales linearly with the computation’s trace length $T$, making them impractical for resource-constrained devices and prohibitively expensive for large-scale tasks. This paper overcomes this barrier by constructing, to our knowledge, the first sublinear-space ZKP prover. The core contribution is using an equivalence that reframes the algebraic process of generating a proof as an instance of the classic Tree Evaluation problem. Leveraging a recent space-efficient tree-evaluation algorithm, we design a streaming prover that recursively assembles the proof without ever materializing the full execution trace. The approach reduces prover memory from $\Theta(T)$ to $O(\sqrt{T})$ (up to lower-order polylogarithmic terms) while preserving identical proof size, verifier time, and transcript/security guarantees of the underlying system. This suggests transformation of verifiable computation from a specialized, server-bound task into one feasible on everyday devices, unlocking new applications in decentralized systems, on-device machine learning, and privacy-preserving technologies.
@@ 1 Introduction
Despite this progress, a fundamental bottleneck persists across practical systems: proof generation requires memory linear in the computation’s trace length. To prove a computation of trace length $T$, existing provers materialize the complete execution trace—i.e., machine state at every row in the algebraic intermediate representation (AIR)—consuming $\Theta(T)$ memory. This is not merely an implementation artifact: the algebraic machinery encodes this full trace into polynomials and enforces global relations (e.g., transition and permutation constraints). The memory cost limits provable scale, blocks deployment on mobile/embedded devices, and inflates costs for web-scale computations.
@@ 1.1 Prior Work and Technical Barriers
Prior attempts to reduce memory have improved constants or parallelization, but have not broken the $\Theta(T)$ barrier. We overcome this limitation by replacing global trace materialization with local, recursive cryptographic operations.
@@ 1.2 Main Contribution: A Sublinear-Space Prover
We establish a precise bridge between complexity theory and applied cryptography. We show that generating ZKP commitments is mathematically equivalent …Applying a recent space-efficient Tree Evaluation algorithm [4] yields, to our knowledge, the first sublinear-space prover within mainstream PCS-based SNARK/Plonkish arithmetizations that preserves verifier cost and (for linear PCSs) the proof/transcript of the baseline prover.
@@
There exists a zero-knowledge proof system in which, for a computation of trace length $T$, the prover operates in $O(\sqrt{T}\cdot\log\log T)$ space (equivalently, $O(\sqrt{T})$ up to a $\log\log$ factor). For linear PCS instantiations (e.g., KZG/IPA), the proof size and verifier complexity are unchanged and the transcript has the same distribution as in the standard prover (and is bit-for-bit identical in non-hiding variants).
@@
    $\mathcal{P}^{\prime}$ uses $O(\sqrt{T}\cdot\log\log T)$ space (equivalently, $O(\sqrt{T})$ up to a $\log\log$ factor).
@@ 1.3 Technical Overview
    Space-efficient evaluation. We evaluate the implicit tree using the Cook–Mertz algorithm [4]. Balancing block size $b_{\text{blk}}$ against the $\log|\mathcal{T}_{C}|\cdot\log\log|\mathcal{T}_{C}|$ stack term yields overall space $O\!\big{(}b_{\text{blk}}+\log|\mathcal{T}_{C}|\cdot\log\log|\mathcal{T}_{C}|\big{)}$ for the traversal plus $O(b_{\text{blk}})$ for local work, optimized at $b_{\text{blk}}=\Theta(\sqrt{T})$, giving $O(\sqrt{T})$ space up to a $\log\log$ factor (Section 5). The prover time is $O(T\log N)\cdot\mathrm{poly}(\lambda)$ for evaluation-basis KZG/IPA with domain size $N\!\geq\!T$ (and $N=\mathrm{poly}(T)$ in typical settings).
@@ 2.1 Notation
We denote the security parameter by $\lambda$. Let $\mathbb{F}$ be a finite field with $|\mathbb{F}|\geq\mathrm{poly}(T)$ so that Schwartz–Zippel error is at most $1/\mathrm{poly}(T)$. Cryptographic parameters are chosen so that the commitment group $\mathcal{C}$ (used by the PCS below) has order at least $2^{\lambda}$. When convenient, we assume $\lambda\geq c\log T$, which implies $1/\mathrm{poly}(T)=2^{-\Omega(\lambda)}$; this relationship is not required but simplifies asymptotic notation.
If $T\mid(|\mathbb{F}|{-}1)$ we fix a primitive $T$-th root of unity $\omega\in\mathbb{F}$ and use the domain $H=\{\omega^{0},\ldots,\omega^{T-1}\}$. Otherwise we choose a domain of size $N\geq T$ either by (i) passing to an extension field $\mathbb{K}\supseteq\mathbb{F}$ that contains a multiplicative subgroup of order $N$, or (ii) using a size-$N$ coset domain inside $\mathbb{F}$ (typically $N$ a power of two), and we pad the trace to length $N$. All polynomial interpolations below are then taken over the chosen domain of size $N$, with degree <N<N; when $N=T$ this specializes to the first case.
All polynomials we manipulate have degree at most $D<N$, yielding Schwartz–Zippel error $D/|\mathbb{F}|\leq 1/\mathrm{poly}(T)$. A function $\nu:\mathbb{N}\to\mathbb{R}$ is negligible if for every positive polynomial $p(\cdot)$, there exists an integer $n_{0}$ such that for all $n>n_{0}$, $\nu(n)<1/p(n)$.
@@ 2.3 Polynomial Commitment Schemes
For non-hiding KZG, commitments are deterministic and (1) holds with $r_{f}=r_{g}=0$. For hiding KZG/IPA variants, $r_{f},r_{g}\!\leftarrow\!\mathbb{F}$ independent uniform implies $r_{f}{+}r_{g}$ is uniform, preserving commitment distributions under linear aggregation.
@@
Suppose the baseline prover commits to each trace/derived polynomial $F_{j}$ in the same basis/SRS fixed above. If $F_{j}=\sum_{t}G_{t,j}$ is the block decomposition and the block commitments $\{\mathsf{Commit}(G_{t,j};r_{t,j})\}_{t}$ are aggregated as in Lemma 2, then for non-hiding PCSs the commitment to $F_{j}$ is bit-for-bit identical to the baseline; for hiding PCSs it has the same distribution.
Quotient construction (coefficient basis).
Let $R$ denote the randomized linear combination of constraint residual polynomials and let $Z_{H}$ be the vanishing polynomial of $H$. We define $Q$ by the polynomial identity
@@
and compute $Q$ in coefficient basis via a streaming backward division by $Z_{H}$, emitting coefficient blocks and committing to them blockwise. By Lemma 2, these block commitments aggregate to the baseline commitment to $Q$ (for linear PCSs), and the approach avoids divide-by-zero at points of $H$.
@@
To open a committed polynomial at a challenge point $\zeta\!\notin\!H$, a Lagrange/evaluation-basis prover can compute $f(\zeta)$ in a single streaming pass using barycentric evaluation, maintaining $O(1)$ accumulators; coefficient-basis provers use a Horner-style pass over streamed coefficients. In both cases, no full polynomial materialization is required, preserving the space bounds.
@@ 3.1 Block-Respecting Decomposition
Any arithmetic circuit $C$ of size $T$ with execution trace matrix $W\in\mathbb{F}^{T\times k}$ naturally yields a $(b_{\text{blk}},\lceil T/b_{\text{blk}}\rceil)$-block-respecting computation with no asymptotic overhead in time or space.
@@ 3.2 The Computation Graph
Crucially for the space-efficient prover, the structure of $G_{C}$ need not be stored. Given the static description of the arithmetic circuit $C$ and an identifier for a node $(m,t)$, its predecessors can be enumerated in $O(\log T)$ space by scanning the local transition template and the circuit’s wiring.
@@
Each computation node $(m,t)$ in $G_{C}$ has indegree at most $\min\{k,\,1{+}r_{\text{reg}}\}$, where $r_{\text{reg}}$ is the cross-register read-degree from the definition above.
@@
We define a function $\mathsf{SimBlock}$ that represents the local computation occurring at each node.
Algorithm 1 SimBlock
0: Boundary vectors $\{B_{0},B_{1},\ldots,B_{r_{\text{reg}}}\}$ from predecessor nodes in block t1t{-}1 (including $B_{0}$ from (m,t1)(m,t{-}1)), time block index $t$, register $m$
0: Output data block $D_{\text{out}}\in\mathbb{F}^{b_{\text{blk}}}$ for register $m$ in time block $t$
@@
2:for each row $i$ in time block $t$ do
3:  $\text{state}\leftarrow\mathsf{F}\big{(}\text{state}\big{)}$ {apply the local transition map induced by the fixed polynomials $P_{1},\ldots,P_{r}$}
4:  $D_{\text{out}}[i]\leftarrow\text{state}[m]$
5:end for
6:return  $D_{\text{out}}$
Lemma 5 (Graph Simulability).
Given the contents of all predecessor nodes of a computation node $(m,t)$, the content of $(m,t)$ can be computed in time $O\!\big{(}(1{+}r_{\text{reg}})\,b_{\text{blk}}\big{)}$ and space $O\!\big{(}b_{\text{blk}}+\deg^{-}(m,t)\big{)}$ using Algorithm 1. In particular, if the AIR has constant read-degree $r_{\text{reg}}=O(1)$ (hence $\deg^{-}(m,t)=O(1)$), this is $O(b_{\text{blk}})$ time and $O(b_{\text{blk}})$ space.
@@
Algorithm 1 takes as input only the predecessor boundary vectors, whose total size is $O(\deg^{-}(m,t))$. Each iteration updates the local state via a constant-arity transition map $\mathsf{F}$ (dependent only on $P_{1},\ldots,P_{r}$) and writes one entry of $D_{\text{out}}$, yielding $b_{\text{blk}}$ iterations and the stated time bound. Storing the predecessor boundaries and the output requires $O(\deg^{-}(m,t)+b_{\text{blk}})$ field elements. ∎
@@ 4.1 From Computation Graphs to Implicit Trees
We never materialize $\mathcal{T}_{C}$ explicitly. The tree view is used to establish distributional equivalence of transcripts; for execution, we evaluate the underlying DAG layer-by-layer so that each node is computed exactly once and retained only until its $O(1)$ consumers read its $\mathsf{aux}$. This schedule achieves the same asymptotic stack bound as the tree traversal (up to an additive $O(\lambda)$ buffer). The height is $h=B=O(T/b_{\text{blk}})$ and the branching factor is $\Delta\leq\min\{k,\,1{+}r_{\text{reg}}\}=O(k)$ (see §3).
@@ 4.2 Tree Functions as Cryptographic Generators
We lift the computation from field elements to cryptographic state. Each node of the computation tree returns (i) a one-hot vector of commitments (nonzero only at its own register coordinate) and (ii) a compact auxiliary state $\mathsf{aux}$ carrying boundary data and running accumulators for global checks (permutation, lookups).
@@
Consequently, concatenating blocks in time order yields the same $Z$ as the monolithic run; commutativity of multiplication implies the same final value for any block schedule consistent with row order.
@@
4: Form $G_{t,m}$ by interpolating regM_vals on $H_{t}$ and padding 0 off $H_{t}$
@@
At each $(m,t)$ we carry forward only coordinate $m$ and never duplicate commitments across registers. A virtual aggregator node $v_{\mathrm{agg}}$ with children $\{(m,B)\}_{m\in[k]}$ reassembles the final $k$-tuple by placing each incoming $m$-coordinate into position $m$. By Lemma 2 and Corollary 1, for linear PCSs the resulting tuple matches (in distribution, and bit-for-bit for non-hiding) the baseline commitments to the global polynomials.

### v2

@@ Abstract
Zero-knowledge proofs allow verification of computations without revealing private information. However, existing systems require memory proportional to the computation size, which has historically limited use in large-scale applications and on mobile and edge devices. We solve this fundamental bottleneck by developing, to our knowledge, the first proof system with sublinear memory requirements for mainstream cryptographic constructions. Our approach processes computations in blocks using a space-efficient tree algorithm, reducing memory from linear scaling to square-root scaling—from $\Theta(T)$ to $O(\sqrt{T}+\log T\log\log T)$ for computation size TT—while maintaining the same proof generation time through a constant number of streaming passes. For widely-used linear polynomial commitment schemes (KZG/IPA), our method produces identical proofs and verification when using the same parameters and hashing only aggregate commitments into the challenge generation, preserving proof size and security. Hash-based systems also achieve square-root memory scaling though with slightly different proof structures. This advance enables zero-knowledge proofs on everyday devices and makes previously infeasible large computations verifiable, fundamentally democratizing access to privacy-preserving computation. Space-efficient zero knowledge proof systems create opportunities to reshape how trust is established in digital systems—from enabling widespread participation in decentralized networks to making verifiable scientific computing practical at unprecedented scales.
@@ 1 Introduction
Despite this progress, a fundamental bottleneck persists across practical systems: proof generation requires memory linear in the computation’s trace length. To prove a computation of trace length $T$, many deployed PCS-based SNARK/Plonkish provers materialize the complete execution trace—i.e., machine state at every row in the algebraic intermediate representation (AIR)—consuming $\Theta(T)$ memory. This is not merely an implementation artifact: the algebraic machinery encodes this full trace into polynomials and enforces global relations (e.g., transition and permutation constraints). The memory cost limits provable scale, blocks deployment on mobile/embedded devices, and inflates costs for web-scale computations.
@@ 1.1 Prior Work and Technical Barriers
Prior attempts to reduce memory have improved constants or parallelization, but have not broken the $\Theta(T)$ barrier in deployed PCS-based systems. We overcome this limitation by replacing global trace materialization with local, recursive cryptographic operations that aggregate to the same global commitments.
@@ 1.2 Main Contribution: A Sublinear-Space Prover
We establish a precise bridge between complexity theory and applied cryptography. We show that the prover’s commitment generation factors into an implicit tree evaluation whose nodes perform local block computations and produce block commitments and auxiliary accumulators. Applying the recent space-efficient Tree Evaluation algorithm [4] yields, to our knowledge, the first sublinear-space prover within mainstream PCS-based SNARK/Plonkish arithmetizations that preserves verifier cost and (for linear PCSs) the proof/transcript of the baseline prover.
@@
There exists a zero-knowledge proof system in which, for a computation of trace length $T$, the prover operates in $O(\sqrt{T}+\log T\log\log T)$ space (equivalently, $O(\sqrt{T})$ up to a $\log\log$ factor). For linear PCS instantiations (e.g., KZG or IPA) and when the transcript exposes only the baseline set of aggregated commitments (no per-block artifacts), the proof size and verifier complexity are unchanged and the transcript has the same distribution as in the standard prover (and is bit-for-bit identical in non-hiding variants).
@@
    $\mathcal{P}^{\prime}$ uses $O(\sqrt{T}+\log T\log\log T)$ space (equivalently, $O(\sqrt{T})$ up to a $\log\log$ factor).
@@ 1.3 Technical Overview
    Space-efficient evaluation. We evaluate the implicit tree using the Cook–Mertz algorithm [4]. Balancing block size $b_{\text{blk}}$ against the $\log|\mathcal{T}_{C}|\cdot\log\log|\mathcal{T}_{C}|$ stack term yields overall space $O\!\big(b_{\text{blk}}+\log|\mathcal{T}_{C}|\cdot\log\log|\mathcal{T}_{C}|\big)$ for the traversal plus $O(b_{\text{blk}})$ for local work, optimized at $b_{\text{blk}}=\Theta(\sqrt{T})$, giving $O(\sqrt{T}+\log T\log\log T)$ space (Section 5). The prover time is $O(T\cdot\mathrm{polylog}\,N)\cdot\mathrm{poly}(\lambda)$ for evaluation-basis KZG/IPA with domain size $N\!\geq\!T$ (and $N=\mathrm{poly}(T)$ in typical settings); for hash-based STARK/FRI the standard $O(T\log T)$ hashing terms apply.
@@ 2.1 Notation
We denote the security parameter by $\lambda$. Let $\mathbb{F}$ be a finite field with $|\mathbb{F}|\geq\mathrm{poly}(T)$ so that Schwartz–Zippel error is at most $1/\mathrm{poly}(T)$. Cryptographic parameters are chosen so that the commitment group $\mathcal{C}$ (used by the PCS below) has order at least $2^{\lambda}$. When convenient, we assume $\lambda\geq c\log T$ for simplifying asymptotic bounds, which implies $1/\mathrm{poly}(T)=2^{-\Omega(\lambda)}$; this relationship is not required but simplifies notation.
If $T\mid(|\mathbb{F}|{-}1)$ we fix a primitive $T$-th root of unity $\omega\in\mathbb{F}$ and use the domain $H=\{\omega^{0},\ldots,\omega^{T-1}\}$ with vanishing polynomial $Z_{H}(X)=X^{T}-1$. Otherwise we choose a domain of size $N\geq T$ either by (i) passing to an extension field $\mathbb{K}\supseteq\mathbb{F}$ that contains a multiplicative subgroup of order $N$, or (ii) using a size-$N$ coset domain inside $\mathbb{F}$ (typically $N$ a power of two), and we pad the trace to length $N$. In both cases, we ensure the vanishing polynomial has the form $Z_{H}(X)=X^{N}-c$ for some constant $c$. All polynomial interpolations below are then taken over the chosen domain of size $N$, with degree <N<N; when $N=T$ this specializes to the first case.
All trace and fixed/selector polynomials we manipulate have degree <N<N. The Plonkish quotient polynomial $Q$ satisfies $\deg Q<cN$ for a small constant $c$ determined by the constraint set (typically $c\in\{2,3,4\}$). Accordingly, we provision the PCS/SRS for a maximum degree $D_{\max}\geq cN$. Schwartz–Zippel error for checks at a random point is then at most $D_{\max}/|\mathbb{F}|\leq 1/\mathrm{poly}(T)$. A function $\nu:\mathbb{N}\to\mathbb{R}$ is negligible if for every positive polynomial $p(\cdot)$, there exists an integer $n_{0}$ such that for all $n>n_{0}$, $\nu(n)<1/p(n)$.
@@ 2.3 Polynomial Commitment Schemes
Linearity is with respect to the fixed commitment basis. For non-hiding KZG, commitments are deterministic and (1) holds with $r_{f}=r_{g}=0$. For hiding KZG/IPA variants, $r_{f},r_{g}\!\leftarrow\!\mathbb{F}$ independent uniform implies $r_{f}{+}r_{g}$ is uniform, preserving commitment distributions under linear aggregation. To achieve bit-identity in hiding PCSs, choose block blinders $r_{1},\ldots,r_{B}$ so $\sum r_{i}=r_{\text{base}}$ (sample B1B{-}1 independently, fix the last as the difference); otherwise, distribution-identity suffices.
@@
Suppose the baseline prover commits to each trace/derived polynomial $F_{j}$ in the same basis/SRS fixed above. If $F_{j}=\sum_{t}G_{t,j}$ is the block decomposition and the block commitments $\{\mathsf{Commit}(G_{t,j};r_{t,j})\}_{t}$ are aggregated as in Lemma 2, then for non-hiding PCSs the commitment to $F_{j}$ is bit-for-bit identical to the baseline; for hiding PCSs it has the same distribution. Critical requirement: transcript identity assumes only the aggregated per-polynomial commitments (not per-block commitments) are hashed into Fiat–Shamir in the same order and encoding as the baseline; internal per-block commitments are computation artifacts and never appear in the transcript.
Quotient construction (coefficient basis, streaming).
Let $R$ denote the randomized linear combination of constraint residual polynomials and let $Z_{H}(X)=X^{N}-c$ be the vanishing polynomial of $H$ (for a subgroup or coset). Define $Q$ by the polynomial identity
@@
and commit to $Q$ in coefficient basis without materializing its full coefficient vector:
(i) Compute the coefficients of $R$ in blocks via a blocked inverse NTT (or any equivalent out-of-core linear transform from evaluations to coefficients), using $O(b_{\text{blk}})$ working memory and $O(N\log N)$ field operations overall.
(ii) As each high-to-low coefficient block of $R$ becomes available, perform the simple backward recurrence for division by $X^{N}-c$; scanning $i=d_{R},\ldots,N$ yields $q_{i-N}\leftarrow r_{i}+c\,q_{i}$, and for $i<N$ consistency enforces $r_{i}=-c\,q_{i}$. Stream the resulting $q$-coefficients directly into the commitment MSM.
By Lemma 2, these block commitments aggregate to the baseline commitment to $Q$ (for linear PCSs), and the approach avoids divide-by-zero at points of $H$.
@@
To open a committed polynomial at a challenge point $\zeta\!\notin\!H$, a Lagrange/evaluation-basis prover can compute $f(\zeta)$ in a single streaming pass using barycentric evaluation, maintaining $O(1)$ accumulators. For a multiplicative subgroup (or coset) with $Z_{H}(X)=X^{N}-c$, we have $Z_{H}^{\prime}(\omega^{i})=N(\omega^{i})^{N-1}$ and weights $w_{i}=1/Z_{H}^{\prime}(\omega^{i})$, so
$f(\zeta)\;=\;\frac{\sum_{i=0}^{N-1}\dfrac{w_{i}\,f(\omega^{i})}{\zeta-\omega^{i}}}{\sum_{i=0}^{N-1}\dfrac{w_{i}}{\zeta-\omega^{i}}},$
which is computable by a single pass over $i=0,\ldots,N{-}1$. Coefficient-basis provers use a Horner-style pass over streamed coefficients. In both cases, no full polynomial materialization is required, preserving the space bounds.
@@ 3.1 Block-Respecting Decomposition
Crucially, each block $t$ needs only boundary values from block t1t{-}1 (plus fixed selectors), making the dependency structure layered and enabling space-efficient evaluation.
@@
Any arithmetic circuit $C$ of size $T$ with execution trace matrix $W\in\mathbb{F}^{T\times k}$ naturally yields a $(b_{\text{blk}},\lceil T/b_{\text{blk}}\rceil)$-block-respecting computation with no asymptotic overhead in time or space, provided the cross-register read-degree $r_{\text{reg}}$ remains bounded.
@@ 3.2 The Computation Graph
Implementation note. We do not materialize $G_{C}$; given the AIR template and a streaming description of $C$, the predecessors of $(m,t)$ can be enumerated in $O(\log T)$ space by scanning local wiring.
@@
Each computation node $(m,t)$ in $G_{C}$ has indegree at most $\min\{k,\,1{+}r_{\text{reg}}\}$, where $r_{\text{reg}}$ is the cross-register read-degree from the definition above. When $k$ or $r_{\text{reg}}$ is not constant, the overall complexity bounds scale accordingly.
@@
Lemma 5 (Layered Schedule).
Each node $(m,t)$ has at most $O(k)$ consumers in layer t+1t{+}1. Thus a layer-by-layer schedule computes each node exactly once and can discard it after its consumers in the next layer have processed it, avoiding recomputation blowup.
We define a block evaluator that emits exactly the row-local fields needed for global checks and the target register’s values.
Algorithm 1 $\mathsf{EvalBlock}(m,t,\text{inputs})$
0: Boundary vectors from predecessors at block t1t{-}1 (including (m,t1)(m,t{-}1)); block index $t$; target register $m$
@@
1: Assemble prior-row inputs at the block boundary
2:for each row $i\in H_{t}$ in order do
3:  Advance the AIR state via the fixed polynomials $P_{1},\ldots,P_{r}$
4:  Append $\text{state}[m]$ to regM_vals
5:  Emit the fixed tuple of row-local fields (e.g., $\{w_{c}(i)\}_{c\in[k]}$ or a fixed compressed form) into local
6:end for
7:return $(\text{regM\_vals},\ \text{local},\ \mathsf{boundary}_{\text{out}})$
Lemma 6 (Graph Simulability, block evaluation).
Given the contents of all predecessor nodes of a computation node $(m,t)$, the content of $(m,t)$ can be computed in time $O\!\big(k\,b_{\text{blk}}\big)$ and space $O\!\big(b_{\text{blk}}+\deg^{-}(m,t)\big)$ using Algorithm 1. In particular, if the AIR has constant register count $k=O(1)$ and read-degree $r_{\text{reg}}=O(1)$ (hence $\deg^{-}(m,t)=O(1)$), this is $O(b_{\text{blk}})$ time and $O(b_{\text{blk}})$ space.
@@
Algorithm 1 takes as input only the predecessor boundary vectors, whose total size is $O(\deg^{-}(m,t))$. Each iteration updates the local state via a constant-arity transition map $\mathsf{F}$ (dependent only on $P_{1},\ldots,P_{r}$), writes one entry of regM_vals, and emits the fixed tuple of row-local fields, yielding $b_{\text{blk}}$ iterations and the stated time bound. Storing the predecessor boundaries and the outputs requires $O(\deg^{-}(m,t)+b_{\text{blk}})$ field elements. ∎
@@ 4.1 From Computation Graphs to Implicit Trees
We never materialize $\mathcal{T}_{C}$ or duplicate DAG nodes: we may traverse as a tree with memoization or process the layered DAG; both realize the same stack bound $O(\log|\mathcal{T}_{C}|\log\log|\mathcal{T}_{C}|)$ up to an additive $O(\lambda)$ buffer while avoiding recomputation blowup. The height is $h=B=O(T/b_{\text{blk}})$ and the branching factor is $\Delta\leq\min\{k,\,1{+}r_{\text{reg}}\}=O(k)$ (see §3).
@@ 4.2 Tree Functions as Cryptographic Generators
We lift the computation from field elements to cryptographic state. Each node of the computation tree returns (i) a one-hot vector of commitments (nonzero only at its own register coordinate) and (ii) a compact auxiliary state $\mathsf{aux}$ carrying boundary data and running accumulators for global checks (permutation, lookups). The node state consists of $O(\lambda)$ group/field elements.
@@
Consequently, blocks must be applied in increasing time order to maintain causality; the same final value results from this unique valid schedule.
Committed accumulator column (permutation).
If the protocol commits to the permutation column $Z$, then during $\mathsf{EvalBlock}$ we also stream the per-row values $Z(i)$ within $H_{t}$ using the recurrence above, producing a vector $\text{z\_vals}\in\mathbb{F}^{|H_{t}|}$. Let $G_{t,Z}$ be the degree-<N<N polynomial equal to $Z(i)$ on $H_{t}$ and 0 elsewhere. We commit to $G_{t,Z}$ and aggregate across $t$ using Lemma 2, yielding the baseline commitment to $Z$ (distribution-identical for linear PCSs).
@@
Committed accumulator column (lookups).
If a lookup accumulator column is committed, we analogously stream its per-row values in each $H_{t}$ to obtain $G_{t,\mathrm{L}}$, commit blockwise, and aggregate via Lemma 2 to the baseline lookup-accumulator commitment.
@@
4:Form $G_{t,m}$: If committing in evaluation basis, use regM_vals on $H_{t}$ directly (padding zeros elsewhere); if committing in coefficient basis, interpolate regM_vals on $H_{t}$ and pad 0 off $H_{t}$
@@
At each $(m,t)$ we carry forward only coordinate $m$ and never duplicate commitments across registers. A virtual aggregator node $v_{\mathrm{agg}}$ with children $\{(m,B)\}_{m\in[k]}$ reassembles the final $k$-tuple by placing each incoming $m$-coordinate into position $m$. By Lemma 2 and Corollary 1, for linear PCSs the resulting tuple matches (in distribution, and bit-for-bit for non-hiding) the baseline commitments to the global polynomials. For transcript identity we require a linear PCS in the same fixed basis/SRS as the baseline, and the transcript must expose only the aggregated commitments; internal per-block commitments are implementation artifacts, not transcript items.